# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ A.3 Computing sequence probabilities in one forward pass
Given that the core algorithm for $(n,p)$-discoverable extraction is so elegantly concise, we include a (simplified) version of the code in this appendix. Our implementation differs from Hayes et al. [60], which we confirm in discussion with the authors of that work. In that work, the authors generated the suffix a token at a time and summed up the per-generated-token conditional $\log$ probabilities. Here, we observe that this is not necessary. We can get the logits for each token in the sequence with just one forward pass through the model, and effectively only one line of code. (See Listing 1, line 28.) For a $50$-token suffix, this results in up to a $50\times$ decrease in compute—an enormous savings. (It is of course less with KV caching enabled during generation, but still nonzero). This also makes computing this metric cheaper than traditional greedy-decoded discoverable extraction [16], which generates the greedy-decoded sequence (i.e., for a $50$-token suffix, there are $50$ forward passes through the model, even if those passes involve caching). Note that this is the entire interaction with the model in our experiments using our probabilistic extraction procedure: for a given sequence, we perform inference on the GPU, get the logits tensor, and move the logits to the CPU for processing. (There are perhaps additional opportunities for improving the efficiency of our implementation that involve KV-caching and streaming logits, rather than performing a giant matmul. We will explore this in the future.) We are not making any changes to the model in our extraction measurements; we are just getting the raw logits from running a given training-data sequence through the model. (As shown in Listing 1, we do this for a batch of sequences at a time, yielding a batch of logits.)The logits tensors we get are of shape [batch_size, sequence_length, vocabulary_size]. This is a three-dimensional tensor:
@@
So, at position [i, j, k] in this tensor, we have the logit (unnormalized $\log$ probability) value for the i-th sequence in the batch, at the j-th position in the token sequence, for the k-th token in the whole token vocabulary. That is, for a given token sequence in the batch (at [i,j]), we have the logit values for all of the next-possible tokens in the whole token vocabulary (a vocabulary_size-length list of numbers, one for each token in the whole vocabulary); these values can be normalized to compute the probability distribution over the next token in the sequence. We then implement our sequence probability computations (i.e., compute $p_{\bm{z}}$) as a post-processing operation on the logits we get from the model. (See Listing 2.) That is, for any given decoding scheme, we can transform the logits—for example, change the shape according to temperature, truncate them to the top-kk—before normalizing them into the $\log$ probability distribution.
@@
This computation is written out on the right side of Equation A.2. The key point here is that, for a given sequence that we are evaluating, all we do (starting at the first index in the suffix) is sum up the conditional $\log$ probabilities of the actual next tokens in the sequence, one at a time (Listing 2, lines 32-40). This is often called teacher forcing.We work with the conditional $\log$ probabilities (adding them together), rather than the conditional probabilities (and multiplying them together), as this is more numerically stable to compute. However, note that these computations are equivalent. For example, consider the sequence "the orange cat" (where, for simplicity, we will assume each word is a single token). When processing "the", the logits reflect the probabilities of the next token. We find the logit value in the distribution that corresponds to "orange" and add the associated conditional $\log$ probability, and then continue on (doing the same for "orange", where we then get the conditional $\log$ probability for the next token being "cat"). We get the conditional $\log$ probability of the whole suffix by adding together the conditional, per-token $\log$ probabilities of the actual suffix. From the suffix $\log$ probability, we can easily obtain the probability $p_{\bm{z}}$: the probability of the model (under the given decoding scheme we are using) generating that suffix conditioned on the prior context (the prefix). Any given sequence should be low probability. That is, for a suffix length of $50$ tokens, this computation is effectively multiplying together $50$ per-token conditional probabilities (by adding together conditional $\log$ probabilities). For a sequence where each token has really high probability, this can still mean the overall sequence has low probability. Consider that each token in a suffix has conditional probability $0.9$. This is really high; it means that a single token in the logit vector (at each position in the suffix) has $90\%$ of the probability mass; for a token vocabulary of size $32,000$, that means the remaining $31,999$ tokens all share the remaining $10\%$. For this suffix, the probability of generating it is $0.9^{50}\approx 0.005$—i.e., a $0.5\%$ chance of generation. This is a relatively small number in the scheme of things; but it is also a really large number in general, if we consider what this means for the underlying token probabilities. In some cases in this paper, we observe sequences that have over $90\%$ probability—an enormous number, especially considering what the underlying computation is. Such unusually high probabilities associated with verbatim suffixes in the training data are indicative of memorization (where it is reasonable to consider $0.5\%$, or even smaller, to be unusually high). As Cooper and Grimmelmann [28] say, “the pattern [the model has learned from the training data] is the memorized training data.” Also note that top-$k$ sampling truncates the logit vector to only the top-$k$ tokens (e.g., for a vocabulary of 32,00032{,}000 tokens, there are 32,00032{,}000 logits, but top-$k$ with k=40k{=}40 decoding will only consider the $40$ highest-valued logits). This means that, if a given sequence has a token at a given position that is not in the top-$k$ of the logit vector, then we will not be able to complete the probability computation. This is what the has_impossible_token check is referring to in Listing 2, lines 36–38. In this case, the given sequence has 0 probability; it is not extractable with any probability.In practice, our code deviates from Listing 2 for efficiency reasons. (The core logic is nevertheless the same as what we show here.) As we process conditional $\log$ probabilities, we save a lot of metadata in order to produce the plots that we include in this paper. For additional speedups (which, ultimately, total around $200\times$ prior work on $(n,p)$-discoverable extraction), as we compute logits on the GPU (Listing 1), we delegate our post-processing code to happen in parallel on CPU.
@@ Appendix B Testing our measurement pipeline
To confirm the accuracy of our implementation of probabilistic discoverable extraction, as well as to benchmark its efficiency in comparison to the implementation in Hayes et al. [60], we re-run the experiments from Hayes et al. [60] for small Pythia models on the Enron dataset. To ensure consistency, we discussed the original run times and sample of Enron with the authors of Hayes et al. [60]. We provide some brief details about these experiments to show due diligence for our implementation in our particular setting. We ran these experiments with both float16 and float32, and report results for float32 to align with Hayes et al. [60]. (However, we note no significant differences between the two.) In Table 1, we show the runtime for our experiments on the 10,00010{,}000, $100$-token sequences Hayes et al. [60] drew from Enron. In Table 2, we report the same metrics as Hayes et al. [60]: the greedy-decoded discoverable extraction rate and (as a point of comparison) the maximum probabilistic rate. We run these experiments on the same $4$ A100s as all of our other probabilistic extraction experiments.
@@
Pythia 1B on Enron subset Hayes et al. [60] (No BOS token) With BOS token
@@
Pythia 2.8B on Enron subset Hayes et al. [60] (No BOS token) With BOS token
@@
Table 2: Greedy discoverable extraction and maximum probabilistic rate for Pythia models on a 10,00010{,}000-sequence subset from Enron. We successfully replicate the exact results in Hayes et al. [60], which did not include the beginning of sequence (BOS) token in the sequences they use for their experiments using Pythia models. We compare to the extraction rates where the BOS token is included at the start of the sequence. Note that the probabilistic rate increases slightly with the presence of the BOS token.
We successfully replicate (exactly) the results of these experiments in Hayes et al. [60]. However, in the process of doing so, we identified a small bug in their original implementation. The Pythia tokenizer is a GPT-2-style tokenizer, and does not add a beginning of sequence (BOS) token by default to the sequences it tokenizes. However, for our setting, it is best practice to include this token at the start of a sequence, as we submit these sequences as inputs for inference to LLMs (that are trained to expect this token at the beginning of the sequence). When we manually prepend this token and re-run the experiments from Hayes et al. [60], we observe elevated extraction rates compared to Hayes et al. [60]. See Table 2. We similarly prepend the BOS token in our Books3 experiments using Pythia models.
@@ Appendix C Experiments on extraction rates
Let $\mathbb{Z}$} be a set of sequences $\bm{z}$} drawn from some (part of a) training dataset. Each $\bm{z}$} is split into a prefix ${\bm{z}}_{1:a}$ and target suffix ${\bm{z}}_{a+1:a+j}$ (e.g., a=j=50a{=}j{=}50 for $100$-token sequences). All extraction rates take the form
@@
where $s({\bm{z}})$ is a success predicate defined according to the chosen criterion. Fix a model $\theta$ and a decoding scheme $\phi$ (e.g., base distribution T=1T{=}1, or top-$k$ with k=40k{=}40), which induces a conditional next-token distribution that can be used to compute suffix probabilities $p_{\bm{z}}\in[0,1]$. For a sequence $\bm{z}$} with prefix ${\bm{z}}_{1:a}$ and suffix ${\bm{z}}_{a+1:a+j}$, define the suffix probability under $\theta$ and $\phi$ as
@@
This is the probability that the model generates the exact target suffix token by token under decoding scheme $\phi$; it is the same definition as Equation A.2 that we use in Section 2 and Appendix A. For traditional, greedy-decoded discoverable extraction [18, 74] (Section 2), $\phi{=}\mathsf{greedy}$ (i.e., T=1,k=1T{=}1,k{=}1) and the success predicate can be written simply as explicit string equality:
@@ C.1 Setup
We perform our average extraction rate experiments with the Books3 dataset. We obtained this dataset from a previous (2022) download of The Pile [47], which is stored on a university cluster and which we use for research purposes on language modeling. The status of Books3 as a research artifact (as well as a training corpus for LLMs) remains unresolved. (See Section 3.) Nevertheless, this dataset remains widely available in the research community; for example, it can be found in the HuggingFace-hosted version of The Pile [42], as well as in various other data repositories [e.g., 113, 3, 14]. It also features as part of a benchmark task in the popular HELM evaluation suite.For this set of experiments, we draw random samples from Books3, performing $3$ separate runs each on 40,00040{,}000 sequences. We discuss how we sample these sets of sequences in Appendix C.2.
@@
From $13$ different models, we try to extract 40,00040{,}000 sequences from Books3 ($3$ times each). We test Llama 1 7B, Llama 1 13B, Llama 1 65B [127], Llama 2 7B, Llama 2 13B, Llama 2 70B [128], Llama 3.1 8B, Llama 3.1 70B [54], Pythia 6.9B, Pythia 12B [9], DeepSeek v1 7B [35], Gemma 2 9B [121], and Phi 4 [2] (a 14B model). We group these models into three size groups: small, medium and large. We use float16 for all models except for Llama 3 and Llama 3.1 models (and Qwen 2.5 models in sliding-window experiments). These models were explicitly trained to work with bfloat16. We know for certain that Pythia and Llama models were trained on Books3, and can conclude from our validity experiments (Section 5 and Appendix E) that DeepSeek v1, Qwen 2.5 and Gemma 2 models were trained on (at least parts of) some books that are also contained in Books3. Phi 4 was deliberately not trained on copyrighted (whole) books, but we find that we can still extract some snippets from copyrighted books.
@@ C.3 High-level takeaways
We defer discussion of validity to Appendix E, where we talk about how to draw conclusions for different models—some for which we know with certainty that Book3 is in their training data, and others for which we can conclude that at least some content that overlaps with Books3 was in their training data.We attempt to use Phi 4 as a reference. Unlike the other models we test, Phi 4 was trained predominantly on synthetic data. Based on the Phi 4 technical report [2], Phi 4 was deliberately not trained on copyrighted (whole) books. As a result, Books3 (the specific corpus) was almost certaintly not included in its training data. Nevertheless, we are sometimes able to extract data that is contained in Books3. These sequences likely found their way into the training data from web-scraped sources that quote these books. We discuss this further in Section 5, Appendix E, and Appendix G. Figure 12 shows average extraction rates grouped roughly by model size. As we expected (and discuss in Section 3), average extraction rates are relatively low. Using the very conservative threshold of $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$, the lowest rate we observe is for Phi 4 ( $0.2\%$), and the highest is for Llama 3.1 70B ($\sim\!0.6\%$). Consistent with prior work, our results show that larger models tend to memorize more than smaller ones [60, 18, 93, 94]. This holds within the same model family. Phi 4 is an exception; despite being a 14B model, its extraction rates on Books3 sequences are comparable to the smaller (7–9B) models we evaluate. Small Llama 3 models memorize on-par with medium Llama 1 and 2 models. In general, later generations of Llama model families tend to memorize more Books3-contained text. For just one example, Llama 3.1 70B memorizes more on average than Llama 2 70B, which memorizes more on average than Llama 1 65B. This pattern does not always hold for specific books. We defer additional, related observations about Llama 3.1 70B to Appendix G.Altogether, it is important to keep in mind that any given sequence that we deemed extracted is, by definition, a member of the training data; however, it is not necessarily a member of the training data because it came from the Books3 corpus. The sequence may be part of another source (or multiple other sources) in the training data. For Pythia and Llama models, the Books3 copy of a sequence may just be one of several duplicate copies in the whole training dataset. The same can be said for models like Gemma 2, where we are not sure if the Books3 corpus was included in the training dataset.) Nevertheless, it is clear from our results that sequences that are contained in Books—but perhaps originate from other sources—are included in the training data (Section 5 & Appendix E).

### v4

@@ A.3 Computing sequence probabilities in one forward pass
Given that the core algorithm for $(n,p)$-discoverable extraction is so elegantly concise, we include a (simplified) version of the code in this appendix. Our implementation differs from hayes2025measuringmemorizationlanguagemodels, which we confirm in discussion with the authors of that work. In that work, the authors generated the suffix a token at a time and summed up the per-generated-token conditional $\log$ probabilities. Here, we observe that this is not necessary. We can get the logits for each token in the sequence with just one forward pass through the model, and effectively only one line of code. (See Listing 1, line 28.) For a $50$-token suffix, this results in up to a $50\times$ decrease in compute—an enormous savings. (It is of course less with KV caching enabled during generation, but still nonzero). This also makes computing this metric cheaper than traditional greedy-decoded discoverable extraction [carlini2021extracting], which generates the greedy-decoded sequence (i.e., for a $50$-token suffix, there are $50$ forward passes through the model, even if those passes involve caching).
Note that this is the entire interaction with the model in our experiments using our probabilistic extraction procedure: for a given sequence, we perform inference on the GPU, get the logits tensor, and move the logits to the CPU for processing. (There are perhaps additional opportunities for improving the efficiency of our implementation that involve KV-caching and streaming logits, rather than performing a giant matmul. We will explore this in the future.) We are not making any changes to the model in our extraction measurements; we are just getting the raw logits from running a given training-data sequence through the model. (As shown in Listing 1, we do this for a batch of sequences at a time, yielding a batch of logits.)
The logits tensors we get are of shape [batch_size, sequence_length, vocabulary_size]. This is a three-dimensional tensor:
@@
So, at position [i, j, k] in this tensor, we have the logit (unnormalized $\log$ probability) value for the i-th sequence in the batch, at the j-th position in the token sequence, for the k-th token in the whole token vocabulary. That is, for a given token sequence in the batch (at [i,j]), we have the logit values for all of the next-possible tokens in the whole token vocabulary (a vocabulary_size-length list of numbers, one for each token in the whole vocabulary); these values can be normalized to compute the probability distribution over the next token in the sequence.
We then implement our sequence probability computations (i.e., compute $p_{\bm{z}}$) as a post-processing operation on the logits we get from the model. (See Listing 2.) That is, for any given decoding scheme, we can transform the logits—for example, change the shape according to temperature, truncate them to the top-kk—before normalizing them into the $\log$ probability distribution.
@@
This computation is written out on the right side of Equation A.2. The key point here is that, for a given sequence that we are evaluating, all we do (starting at the first index in the suffix) is sum up the conditional $\log$ probabilities of the actual next tokens in the sequence, one at a time (Listing 2, lines 32-40). This is often called teacher forcing.
We work with the conditional $\log$ probabilities (adding them together), rather than the conditional probabilities (and multiplying them together), as this is more numerically stable to compute. However, note that these computations are equivalent. For example, consider the sequence "the orange cat" (where, for simplicity, we will assume each word is a single token). When processing "the", the logits reflect the probabilities of the next token. We find the logit value in the distribution that corresponds to "orange" and add the associated conditional $\log$ probability, and then continue on (doing the same for "orange", where we then get the conditional $\log$ probability for the next token being "cat"). We get the conditional $\log$ probability of the whole suffix by adding together the conditional, per-token $\log$ probabilities of the actual suffix. From the suffix $\log$ probability, we can easily obtain the probability $p_{\bm{z}}$: the probability of the model (under the given decoding scheme we are using) generating that suffix conditioned on the prior context (the prefix).
Any given sequence should be low probability. That is, for a suffix length of $50$ tokens, this computation is effectively multiplying together $50$ per-token conditional probabilities (by adding together conditional $\log$ probabilities). For a sequence where each token has really high probability, this can still mean the overall sequence has low probability.
Consider that each token in a suffix has conditional probability $0.9$. This is really high; it means that a single token in the logit vector (at each position in the suffix) has $90\%$ of the probability mass; for a token vocabulary of size $32,000$, that means the remaining $31,999$ tokens all share the remaining $10\%$. For this suffix, the probability of generating it is $0.9^{50}\approx 0.005$—i.e., a $0.5\%$ chance of generation. This is a relatively small number in the scheme of things; but it is also a really large number in general, if we consider what this means for the underlying token probabilities. In some cases in this paper, we observe sequences that have over $90\%$ probability—an enormous number, especially considering what the underlying computation is.
Such unusually high probabilities associated with verbatim suffixes in the training data are indicative of memorization (where it is reasonable to consider $0.5\%$, or even smaller, to be unusually high). As cooper2024files say, “the pattern [the model has learned from the training data] is the memorized training data.”
Also note that top-$k$ sampling truncates the logit vector to only the top-$k$ tokens (e.g., for a vocabulary of 32,00032{,}000 tokens, there are 32,00032{,}000 logits, but top-$k$ with k=40k{=}40 decoding will only consider the $40$ highest-valued logits). This means that, if a given sequence has a token at a given position that is not in the top-$k$ of the logit vector, then we will not be able to complete the probability computation. This is what the has_impossible_token check is referring to in Listing 2, lines 36–38. In this case, the given sequence has 0 probability; it is not extractable with any probability.
In practice, our code deviates from Listing 2 for efficiency reasons. (The core logic is nevertheless the same as what we show here.) As we process conditional $\log$ probabilities, we save a lot of metadata in order to produce the plots that we include in this paper. For additional speedups (which, ultimately, total around $200\times$ prior work on $(n,p)$-discoverable extraction), as we compute logits on the GPU (Listing 1), we delegate our post-processing code to happen in parallel on CPU.
@@ Appendix B Testing our measurement pipeline
To confirm the accuracy of our implementation of probabilistic discoverable extraction, as well as to benchmark its efficiency in comparison to the implementation in hayes2025measuringmemorizationlanguagemodels, we re-run the experiments from hayes2025measuringmemorizationlanguagemodels for small Pythia models on the Enron dataset. To ensure consistency, we discussed the original run times and sample of Enron with the authors of hayes2025measuringmemorizationlanguagemodels. We provide some brief details about these experiments to show due diligence for our implementation in our particular setting.
We ran these experiments with both float16 and float32, and report results for float32 to align with hayes2025measuringmemorizationlanguagemodels. (However, we note no significant differences between the two.) In Table 1, we show the runtime for our experiments on the 10,00010{,}000, $100$-token sequences hayes2025measuringmemorizationlanguagemodels drew from Enron. In Table 2, we report the same metrics as hayes2025measuringmemorizationlanguagemodels: the greedy-decoded discoverable extraction rate and (as a point of comparison) the maximum probabilistic rate. We run these experiments on the same $4$ A100s as all of our other probabilistic extraction experiments.
@@
Pythia 1B on Enron subset hayes2025measuringmemorizationlanguagemodels (No BOS token) With BOS token
@@
Pythia 2.8B on Enron subset hayes2025measuringmemorizationlanguagemodels (No BOS token) With BOS token
@@
Table 2: Greedy discoverable extraction and maximum probabilistic rate for Pythia models on a 10,00010{,}000-sequence subset from Enron. We successfully replicate the exact results in hayes2025measuringmemorizationlanguagemodels, which did not include the beginning of sequence (BOS) token in the sequences they use for their experiments using Pythia models. We compare to the extraction rates where the BOS token is included at the start of the sequence. Note that the probabilistic rate increases slightly with the presence of the BOS token.
We successfully replicate (exactly) the results of these experiments in hayes2025measuringmemorizationlanguagemodels. However, in the process of doing so, we identified a small bug in their original implementation. The Pythia tokenizer is a GPT-2-style tokenizer, and does not add a beginning of sequence (BOS) token by default to the sequences it tokenizes. However, for our setting, it is best practice to include this token at the start of a sequence, as we submit these sequences as inputs for inference to LLMs (that are trained to expect this token at the beginning of the sequence). When we manually prepend this token and re-run the experiments from hayes2025measuringmemorizationlanguagemodels, we observe elevated extraction rates compared to hayes2025measuringmemorizationlanguagemodels. See Table 2. We similarly prepend the BOS token in our Books3 experiments using Pythia models.
@@ Appendix C Experiments on extraction rates
Let $\mathbb{Z}$} be a set of sequences $\bm{z}$} drawn from some (part of a) training dataset. Each $\bm{z}$} is split into a prefix ${\bm{z}}_{1:a}$ and target suffix ${\bm{z}}_{a+1:a+j}$ (e.g., a=j=50a{=}j{=}50 for $100$-token sequences).
All extraction rates take the form
@@
where $s({\bm{z}})$ is a success predicate defined according to the chosen criterion.
Fix a model $\theta$ and a decoding scheme $\phi$ (e.g., base distribution T=1T{=}1, or top-$k$ with k=40k{=}40), which induces a conditional next-token distribution that can be used to compute suffix probabilities $p_{\bm{z}}\in[0,1]$. For a sequence $\bm{z}$} with prefix ${\bm{z}}_{1:a}$ and suffix ${\bm{z}}_{a+1:a+j}$, define the suffix probability under $\theta$ and $\phi$ as
@@
This is the probability that the model generates the exact target suffix token by token under decoding scheme $\phi$; it is the same definition as Equation A.2 that we use in Section 2 and Appendix A.
For traditional, greedy-decoded discoverable extraction [carlini2023quantifying, lee2022dedup] (Section 2), $\phi{=}\mathsf{greedy}$ (i.e., T=1,k=1T{=}1,k{=}1) and the success predicate can be written simply as explicit string equality:
@@ C.1 Setup
We perform our average extraction rate experiments with the Books3 dataset. We obtained this dataset from a previous (2022) download of The Pile [gao2020pile], which is stored on a university cluster and which we use for research purposes on language modeling.
The status of Books3 as a research artifact (as well as a training corpus for LLMs) remains unresolved. (See Section 3.) Nevertheless, this dataset remains widely available in the research community; for example, it can be found in the HuggingFace-hosted version of The Pile [hfpilededup], as well as in various other data repositories [e.g., books3a, books3b, books3c]. It also features as part of a benchmark task in the popular HELM evaluation suite.
For this set of experiments, we draw random samples from Books3, performing $3$ separate runs each on 40,00040{,}000 sequences. We discuss how we sample these sets of sequences in Appendix C.2.
@@
From $13$ different models, we try to extract 40,00040{,}000 sequences from Books3 ($3$ times each). We test Llama 1 7B, Llama 1 13B, Llama 1 65B [touvron2023llamaopenefficientfoundation], Llama 2 7B, Llama 2 13B, Llama 2 70B [llama2], Llama 3.1 8B, Llama 3.1 70B [llama3], Pythia 6.9B, Pythia 12B [biderman2023pythia], DeepSeek v1 7B [deepseekv1], Gemma 2 9B [gemma2], and Phi 4 [phi4] (a 14B model). We group these models into three size groups: small, medium and large.
We use float16 for all models except for Llama 3 and Llama 3.1 models (and Qwen 2.5 models in sliding-window experiments). These models were explicitly trained to work with bfloat16.
We know for certain that Pythia and Llama models were trained on Books3, and can conclude from our validity experiments (Section 5 and Appendix E) that DeepSeek v1, Qwen 2.5 and Gemma 2 models were trained on (at least parts of) some books that are also contained in Books3. Phi 4 was deliberately not trained on copyrighted (whole) books, but we find that we can still extract some snippets from copyrighted books.
@@ C.3 High-level takeaways
We defer discussion of validity to Appendix E, where we talk about how to draw conclusions for different models—some for which we know with certainty that Book3 is in their training data, and others for which we can conclude that at least some content that overlaps with Books3 was in their training data.
We attempt to use Phi 4 as a reference. Unlike the other models we test, Phi 4 was trained predominantly on synthetic data. Based on the Phi 4 technical report [phi4], Phi 4 was deliberately not trained on copyrighted (whole) books. As a result, Books3 (the specific corpus) was almost certaintly not included in its training data. Nevertheless, we are sometimes able to extract data that is contained in Books3. These sequences likely found their way into the training data from web-scraped sources that quote these books. We discuss this further in Section 5, Appendix E, and Appendix G.
Figure 12 shows average extraction rates grouped roughly by model size. As we expected (and discuss in Section 3), average extraction rates are relatively low. Using the very conservative threshold of $p_{\bm{z}}\geq\tau_{\text{min}}=0.1\%$, the lowest rate we observe is for Phi 4 ( $0.2\%$), and the highest is for Llama 3.1 70B ($\sim\!0.6\%$). Consistent with prior work, our results show that larger models tend to memorize more than smaller ones [hayes2025measuringmemorizationlanguagemodels, carlini2023quantifying, nasr2023scalable, nasr2025scalable]. This holds within the same model family. Phi 4 is an exception; despite being a 14B model, its extraction rates on Books3 sequences are comparable to the smaller (7–9B) models we evaluate. Small Llama 3 models memorize on-par with medium Llama 1 and 2 models. In general, later generations of Llama model families tend to memorize more Books3-contained text. For just one example, Llama 3.1 70B memorizes more on average than Llama 2 70B, which memorizes more on average than Llama 1 65B. This pattern does not always hold for specific books. We defer additional, related observations about Llama 3.1 70B to Appendix G.
Altogether, it is important to keep in mind that any given sequence that we deemed extracted is, by definition, a member of the training data; however, it is not necessarily a member of the training data because it came from the Books3 corpus. The sequence may be part of another source (or multiple other sources) in the training data. For Pythia and Llama models, the Books3 copy of a sequence may just be one of several duplicate copies in the whole training dataset. The same can be said for models like Gemma 2, where we are not sure if the Books3 corpus was included in the training dataset.) Nevertheless, it is clear from our results that sequences that are contained in Books—but perhaps originate from other sources—are included in the training data (Section 5 & Appendix E).